Solve the Cauchys Linear Differential Equations :

x^2.d^2y/dx^2 + x.dy/dx - y = x^3.e^4


1
Expert's answer
2021-10-11T01:27:46-0400

The homogeneous differential equation


"x^2y''+xy'-y=0"

Substitute "y=x^\lambda"


"x^2\lambda(\lambda-1)x^{\lambda-2}+x\lambda x^{\lambda-1}-x^\lambda=0"

"\lambda^2-1=0"

"\lambda_1=1, \lambda_2=-1"

The general solution to the homogeneous differential equation is


"y_h=c_1x+\dfrac{c_2}{x}"

Find the particular solution to the non homogeneous differential equation


"y_p=Ae^4x^3"

"y_p'=3Ae^4x^2"

"y_p''=6Ae^4x"

Substitute


"x^2(6Ae^4x)+x(3Ae^4x^2)-Ae^4x^3=e^4x^3"

"A=\dfrac{1}{8}"

The particularsolution to the non homogeneous differential equation is


"y_p=\dfrac{1}{8}e^4x^3"


The general solution to the homogeneous differential equation is


"y=y_h+y_p"


"y(x)=c_1x+\dfrac{c_2}{x}+\dfrac{1}{8}e^4x^3"